an object moving in simple harmonic motion has an amplitude of 0.02m and a frequency of 20hz. find the acceleration and velocity

1
Expert's answer
2022-02-28T10:40:47-0500

The equation of SHM

"x=A\cos(2\pi ft)=0.02\cos(40\pi t)"

The velocity

"v=x'=-0.8\pi\sin(40\pi t)"

The acceleration

"a=v'=-32\pi^2\cos(40\pi t)"